# xcolorIncompatible color definition

### v1

@@ Abstract
Current simulators lack the ability to accurately model integrated environments that encompass sea, air, and land. To address this gap, we introduce Aerial-Aquatic Manipulators (AAMs) in SEa, Air, and Land Simulator (SEALS), a comprehensive and photorealistic simulator designed for AAMs to operate and learn in these diverse environments. The development of AAM-SEALS tackles several significant challenges, including the creation of integrated controllers for flying, swimming, and manipulation, and the high-fidelity simulation of aerial dynamics and hydrodynamics leveraging particle physics. Our evaluation demonstrates smooth operation and photorealistic transitions across air, water, and their interfaces. We quantitatively validate the fidelity of particle-based hydrodynamics by comparing position-tracking errors across real-world and simulated systems. AAM-SEALS promises to benefit a broad range of robotics communities, including robot learning, aerial robotics, underwater robotics, mobile manipulation, and robotic simulators. We will open-source our code and data to foster the advancement of research in these fields. Please access our project website at: https://aam-seals.github.io/aam-seals-v1/
@@ I INTRODUCTION
Application: We introduced a novel cross-medium manipulation task involving the search and capture of moving aquatic animals such as crabs. We showcased the effectiveness of our AAM and SEALS by teleoperating the AAM to complete these challenging tasks, resulting in a unique dataset with fully controllable crab models and demonstration trajectories for AAM crab-catching. Additionally, we applied cutting-edge visual reinforcement learning to further enhance AAM performance.
@@ III-A AAM Dynamics Modeling
Figure 1: Schematic of a representative Aerial Aquatic Manipulator.
@@ IV-C Simulation Realism
Realistic Air-Water Transition:
Our SEALS system uses particle-based hydrodynamics to achieve highly realistic air-water transitions, capturing both dynamic interactions and detailed rendering. As discussed in Appendix B, this approach simulates cohesion and surface tension, allowing realistic interactions between fluid and solid surfaces. Grounded in solid theoretical principles, we observe that particle-based hydrodynamics in SEALS effectively simulate water splashes when the AAM impacts the water and damping effects as it transitions from air to water, while the damping effect is demonstrated by deactivating the AAM’s thrusters and allowing it to descend into the water under gravity. This effect causes a sudden change in acceleration as the AAM enters the water. Due to disturbances in the surface of natural water, such as wind-induced waves, the AAM loses the balance it maintains in the air once it submerges. For a detailed demonstration, please refer to the accompanying video: https://shorturl.at/XG9Cr
@@
Light Attenuation:
Light attenuation in water refers to the gradual reduction in the intensity of light as it travels through water. This phenomenon occurs as a result of the absorption and scattering of light by water molecules and suspended particles. Different wavelengths of light are absorbed at varying rates, with longer wavelengths like red being absorbed more quickly, while shorter wavelengths like blue penetrate deeper. This results in objects appearing bluer and more blurred as they move farther away from the observer.
@@
In addition to absorption, light scattering, often observed as caustics, occurs when light rays bend and disperse as they pass through varying densities in the water, creating intricate patterns of light and shadow on underwater surfaces. As a result of these effects, objects appear less bright and can even change color as they move further away from the light source in underwater environments.
@@
Our SEALS system supports these features to further enhance realism beyond what is offered by Isaac Sim. As a result, objects appear less bright and can even change color as they move further away from the light source in underwater environments. Our SEALS supports such features to further enhance realism over Isaac Sim. Fig. 5 shows the control of light intensity with respect to water depth, while Fig. 6 shows the light absorption effects and the light scattering (caustics) effects.
Realistic Wave-Drone Interaction:
In the Fig. 4 and this demo video: https://shorturl.at/aT8yf, we showcase an AAM in free fall that is unexpectedly struck by an ocean wave. The sudden shifts in acceleration caused by the wave impact are quantitatively illustrated in Fig. 12.
Realistic Aquatic Animals:
While most photorealistic simulators focus primarily on sensory realism, our SEALS system is the first to also emphasize the realistic simulation of environmental animal behavior. We have meticulously developed detailed meshes and kinematic models of aquatic animals, such as crabs and sea spiders, and equipped them with controllers that facilitate robot learning for developing control policies, as shown in Fig. 7 and 8.
Figure 7: Simulated crab walking slowly on the sea floor.
Figure 8: Simulated sea spider captured by our Aerial-Aquatic Manipulator.
@@ IV-E Control Interfaces to Robots
IV-F Dataset for Aquatic Animal Search and Capture by AAMs
The dataset includes fully controllable models of a crab and a sea spider, as detailed in Sec. IV-C, along with expert-teleoperated demonstrations of an exemplary AAM performing search-and-capture tasks for aquatic animals. These demonstrations enable reinforcement learning from demonstrations (RLfD), a method that combines the efficiency of imitation learning with the exploratory capabilities of reinforcement learning, as shown in prior works [44, 45]. RLfD typically requires as few as 10 demonstrations to achieve effective learning. The AAM was teleoperated using a joystick. We will release the dataset and corresponding AAM trajectories at: https://shorturl.at/pzski.
@@ V-A Evaluating the Realism of Particle-based Hydrodynamics
In SEALS, we disabled the AAM’s thrusters and allowed it to fall freely from the air into the water. We recorded the changes in the AAM’s acceleration over time, as shown in Fig. 10, which clearly demonstrates a sudden shift in acceleration at the moment the AAM enters the water, caused by water damping effects.
Figure 10: Plot of acceleration along z-direction over time while a simulated AAM falls freely from the air into the water.
For the real-world experiment, we equipped a 3D-printed AAM with an IMU sensor (WitMotion WT901BLECL, 10 Hz output) and dropped it into a 25-foot-deep water tank. The resulting acceleration-over-time plot is shown in Fig.11. Both Fig.10 and Fig. 11 exhibit similar patterns during the air-water transition, with sharp acceleration changes when the object impacts the water surface and the floor. During the sinking phase, the absolute accelerations in both the real-world and simulated scenarios stabilize around 10, though the low output frequency of the real-world IMU may have missed some acceleration peaks.
Figure 11: Left: Our 3D-printed AAM. Middle: The 25-foot-deep water tank used in our experiments. Right: Acceleration plot along the z-direction over time as a real-world AAM freely falls from air into water.
@@ V-B Evaluating the Control of AAM in SEALS Simulator
We evaluate the AAM’s control performance by analyzing position-tracking error across multiple tasks: Task 1 involves hovering control with arm movements to evaluate the robustness of the system to changes in the center of gravity, while Task 2 (to be completed) focuses on following a circular trajectory that transitions between air and water, testing the control architecture’s cross-medium effectiveness.
The position-tracking results, shown in Fig. 13, demonstrate precise control while being robust to the changes of Center of Gravity: the X-axis remains within ±$\pm 0.015m$, the Y-axis within ±$\pm 0.003m$, and the Z-axis within ±$\pm 0.2m$. The demo video can be visualized at: https://aam-seals.github.io/aam-seals-v1/media/videos/hovering_demo1.mp4
V-C Visual Reinforcement Learning Evaluation
We applied the Soft Actor-Critic algorithm [46], a robust reinforcement learning (RL) method for continuous control, to train our AAM in the SEALS environment to reach objects such as crabs in Fig. 9, using distance-based rewards. The observation space included two 128 x 128 RGB images from the front and bottom cameras on the AAM (resulting in six channels in total) and the global poses of the AAM and the target object. The cumulative average rewards over the past 100 episodes, shown in Fig. 8, indicate that the reinforcement learning converges. For more detailed robot learning results, please refer to Appendix C.
@@ VI Conclusion
Our work makes significant contributions to the field of robotics by introducing a new class of robots, Aerial-Aquatic Manipulators (AAMs), and developing the first high-fidelity simulator that integrates sea, air, and land environments. The unique benefits of this work include enabling AAMs to seamlessly transition between different environments and perform complex tasks that require multi-terrain capabilities.
Limitations: While our research presents a robust simulation environment, we acknowledge the limitation of not fully verifying the Sim2Real transfer of AAM systems. Developing a fully functional physical AAM involves numerous challenges, such as seamlessly integrating electronics and on-board GPUs, achieving waterproofing to at least IP68 standards, designing lightweight and robust structures, and ensuring battery safety – all of which could constitute a significant standalone research effort. As an initial step, we conducted experiments with 3D-printed AAMs, including free-fall tests into water, which validated our simulator’s hydrodynamic accuracy by demonstrating realistic acceleration patterns during air-to-water transitions, closely mirroring real-world observations. While these validations miss the full operational complexity of cross-medium tasks, it’s notable that many highly regarded works that develop complex photorealistic simulators, such as [47, 48, 49, 50, 31, 51], also do not include physical robot experiments due to the significant challenges involved. By comparison, AAM-SEALS makes a meaningful contribution by not only providing a state-of-the-art simulation environment but also offering a novel class of robots with carefully designed morphology, kinematics, dynamics, and control systems. This highlights the inherent chicken-and-egg problem, where physical prototypes and simulators evolve iteratively to refine design and functionality. AAM-SEALS initializes this interplay by providing a state-of-the-art simulation environment and introducing the novel class of robots. We believe AAM-SEALS opens a broad avenue for future research efforts:
  • Supporting the Development of Physical AAMs: Leveraging AAM-SEALS to aid in the development of physical AAMs is another ongoing project. This simulator will provide critical insights and validation before constructing physical prototypes.
  • Benchmarking the AAM-SEALS: This work evaluates cutting-edge reinforcement learning techniques, e.g. RLfD, on challenging cross-medium navigation and crab capturing tasks. While those tasks are practically valuable, future work will benchmark paper based on the fundamentals in this work, introducing more valuable tasks depicted in Fig. LABEL:fig:intro such as undersea search and rescue, efficient underwater debries cleaning, on a broader range of cutting-edge reinforcement learning techniques.
@@
ACKNOWLEDGMENT
We would like to thank Marcelo Jacinto, the author of the Pegasus simulator [23], for his insightful debugging advice during a meeting with our team. Our appreciation also goes to Charlie Hanner for providing essential diving support for our real-world experiment, which measured acceleration changes over time as objects fell into water. Additionally, we are grateful to Isha Hemant Surjuse and Ian Fuller for their invaluable guidance and support in 3D-printing the Aerial-Aquatic Manipulator, which was instrumental to the success of this project. Finally, we thank Guanqun Luo for her invaluable contributions to the creation of Fig. LABEL:fig:intro.
@@ Appendix A Potential Questions and Our Answers
While our work does not include experiments on physical AAMs, this reflects the significant challenges involved in developing functional prototypes, such as electronics design, achieving waterproofing, lightweight structures, and battery safety. Instead, we validated our simulator’s hydrodynamic accuracy through experiments with 3D-printed models, showing realistic acceleration patterns during air-to-water transitions. These preliminary tests, however, do not fully capture the complexity of cross-medium tasks. Similar limitations are common in other advanced simulators (e.g., [47, 48, 49, 50, 31, 51]), which also prioritize simulation development over physical experiments due to practical constraints. Our simulation environment serves as a foundational tool to guide future physical AAM development.
@@ Appendix C Additional Reinforcement Learning Experiments
C-A Training Results
The reinforcement learning results are presented in Fig. 15, illustrating the mean of episode cumulative rewards (ep_rew_mean) over steps, the success rate over steps, and the mean of episode length (ep_len_mean) over steps.
Figure 15: Reinforcement Learning Results
C-B Reinforcement Learning Hyperparameters
@@
replay buffer size 1,000,000
@@
learning starts 100

### v3

@@ Abstract
Current mobile manipulators and high-fidelity simulators lack the ability to seamlessly operate and simulate across integrated environments spanning sea, air, and land. To address this gap, we introduce Aerial-Aquatic Manipulators (AAMs) in SEa, Air, and Land Simulator (SEALS), a comprehensive and photorealistic simulator designed for AAMs to operate and learn in these diverse environments. The development of AAM-SEALS tackles several significant challenges, including the creation of integrated controllers for flying, swimming, and manipulation, and the high-fidelity simulation of aerial dynamics and hydrodynamics leveraging particle-based hydrodynamics. Our evaluation demonstrates smooth operation and photorealistic transitions across air, water, and their interfaces. We quantitatively validate the fidelity of particle-based hydrodynamics by comparing position-tracking errors across real-world and simulated systems. AAM-SEALS benefits a broad range of robotics communities, including robot learning, aerial robotics, underwater robotics, mobile manipulation, and robotic simulators. We will open-source our code and data to foster the advancement of research in these fields. The overview video is available at https://youtu.be/MbqIIrYvR78. Visit our project website at https://aam-seals-v1.umd.edu for more details.
**footnotetext: The first four authors contributed equally, and are listed in reverse alphabetical order.
@@ I INTRODUCTION
Application: We introduced a novel aerial-aquatic manipulation challenge centered on searching for and capturing moving aquatic animals, such as crabs. To demonstrate the effectiveness of our AAM and SEALS, we developed a unique dataset featuring fully controllable aquatic animal models and specialized tools for collecting demonstration trajectories for AAM-based animal capture. Additionally, we implemented state-of-the-art reinforcement learning from demonstrations (RLfD) and visual reinforcement learning to facilitate future research in aerial-aquatic manipulation.
@@ III-A AAM Dynamics Modeling
Figure 1: Schematic of a representative Aerial Aquatic Manipulator.
@@ IV-C Simulation Realism
IV-C1 Realistic Air-Water Transition
Our SEALS system uses particle-based hydrodynamics to achieve highly realistic air-water transitions, capturing both dynamic interactions and detailed rendering. As discussed in Appendix B, this approach simulates cohesion and surface tension, allowing realistic interactions between fluid and solid surfaces. Grounded in solid theoretical principles, we observe that particle-based hydrodynamics in SEALS effectively simulate water splashes when the AAM impacts the water and damping effects as it transitions from air to water, while the damping effect is demonstrated by deactivating the AAM’s thrusters and allowing it to descend into the water under gravity. This effect causes a sudden change in acceleration as the AAM enters the water. Due to disturbances in the surface of natural water, such as wind-induced waves, the AAM loses the balance it maintains in the air once it submerges.
IV-C2 Realistic Wave-Drone Interaction
In the Fig. 4, we showcase an AAM in free fall that is unexpectedly struck by an ocean wave. The sudden shifts in acceleration caused by the wave impact are quantitatively illustrated in Fig. 10.
IV-C3 Realistic Underwater Observation
In underwater environments, light attenuation significantly impacts visibility due to absorption and scattering by water molecules and suspended particles. As depth increases, light intensity diminishes, leading to reduced visibility, as shown in Fig.5.
@@ IV-C3 Realistic Underwater Observation
Additionally, different wavelengths are absorbed at varying rates; longer wavelengths like red are absorbed more quickly, while shorter wavelengths like blue penetrate deeper, causing distant objects to appear bluer and more blurred. This depth-dependent color shift and blurriness, illustrated in Fig. 6 (middle), affect how objects are perceived underwater. Furthermore, light scattering, often observed as caustics, occurs when light rays bend and disperse through varying water densities, creating intricate patterns of light and shadow on submerged surfaces, as shown in Fig. 6 (right).
@@
IV-C4 Realistic Aquatic Animals
While most photorealistic simulators focus primarily on sensory realism, our SEALS system is the first to also emphasize the realistic simulation of environmental animal behavior. We have meticulously developed detailed meshes and kinematic models of aquatic animals, such as crabs and sea spiders, and equipped them with controllers that facilitate robot learning for developing control policies, as shown in Fig. 7.
Figure 7: Top: Simulated crab walking slowly on the sea floor; Bottom: Simulated sea spider captured by our Aerial-Aquatic Manipulator.
@@ IV-E Control Interfaces to Robots
IV-F Aquatic Animal Search and Capture Challenge for AAMs
Building upon the capabilities of our SEALS simulator, which features fully controllable models of a crab and a sea spider (as detailed in Sec. IV-C4), we present a novel challenge: the search and capture of aquatic animals. This task demands intricate tacit strategies and the flexibility to adapt learned skills to unforeseen scenarios.
To facilitate the transfer of these implicit strategies to robots, we employ teleoperation of our AAM using a joystick to gather demonstrations for search-and-capture tasks. We advocate for the use of reinforcement learning from demonstrations (RLfD), a method that merges the efficiency of acquiring expert knowledge through imitation learning with the exploratory strengths of reinforcement learning, as shown in prior works [32, 52]. Typically, RLfD requires as few as 10 demonstrations to achieve effective learning. Our demonstrations are available at this link: https://tinyurl.com/5b8sm9fc.
IV-G Reinforcement Learning from Demonstrations (RLfD)
Building on the control interfaces explained in Sec. IV-E, we implement our Reinforcement Learning from Demonstrations (RLfD) framework. We model the environment as a Markov Decision Process (MDP) defined by the tuple $(S,A,P,R,\gamma)$, where $S$ denotes the state space, $A$ the action space, $P(s^{\prime}|s,a)$ the state transition probability, $R(s,a)$ the reward function, and $\gamma$ the discount factor. Our RLfD implementation builds upon Soft Actor-Critic from Demonstrations (SACfD) implemented in [52], which integrates expert demonstrations into Soft Actor-Critic (SAC) [11] to accelerate policy learning while maintaining exploration. SAC is a state-of-the-art RL algorithm for continuous control tasks, which optimizes policies by maximizing both the expected cumulative reward and an entropy term to encourage exploration, our approach integrates expert demonstrations to expedite learning. The objective function in SAC is given by:
$J(\pi)=\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}\left(R(s_{t},a_{t})+\alpha\mathcal{H}(\pi(\cdot|s_{t}))\right)\right]$ (9)
where $\alpha$ is a temperature parameter that balances the reward and entropy terms, and $\mathcal{H}(\pi(\cdot|s_{t}))$ represents the entropy of the policy at state $s_{t}$. To effectively incorporate demonstrations, we employ several strategies:
  • Pretraining: SACfD initializes the agent’s policy and value functions by pretraining on the demonstration data, which provides a strong starting point before interacting with the environment.
  • Prioritized Experience Replay: SACfD utilizes a prioritized replay buffer where demonstration experiences are assigned higher priorities, ensuring that the agent samples these informative transitions more frequently during training.
  • N-Step Returns: To better learn longer-term dependencies from demonstrations and provide more informative updates, SACfD incorporates n-step returns in the learning process, which has the following form:
    $G_{t}=\sum_{k=0}^{N-1}\gamma^{k}R_{t+k}+\gamma^{N}V(s_{t+N})$
    where $V(s_{t+N})$ is the estimated value of the state after $N$ steps. $R_{t+k}$ is the reward at step $t+k$.
Figure 9: Left: A simulated AAM and its acceleration along the z-axis over time as it falls from air into water. Middle: The 3D-printed AAM and the water tank used in real-world experiments. Right: Acceleration along the z-axis over time for the real-world AAM falling into water.
@@ V-A Evaluating the Realism of Particle-based Hydrodynamics
In SEALS, we disabled the AAM’s thrusters and allowed it to fall freely from the air into the water. For the real-world experiment, we equipped a 3D-printed AAM with an IMU sensor (WitMotion WT901BLECL, 50 Hz output) and dropped it into a water tank. We recorded the changes in the AAM’s acceleration over time for both simulated and real world experiments, as shown in Fig. 9. The results exhibit similar patterns in both cases: the acceleration initially starts at gravitational acceleration, then rapidly decreases to near zero due to air resistance. Upon water entry, a sharp fluctuation occurs due to water damping effects, followed by another fluctuation when the AAM hits the ground.
@@ V-B Evaluating the Control of AAM in SEALS Simulator
We evaluate the AAM’s control performance by analyzing position-tracking error across multiple tasks: Task 1 involves hovering control with arm movements to evaluate the robustness of the system to changes in the center of gravity, while Task 2 focuses on following an oval trajectory that transitions between air and water, testing the control architecture’s cross-medium effectiveness.
The position-tracking results of hovering task, shown in Fig. 11, demonstrate precise control while being robust to the changes of Center of Gravity: the X-axis remains within ±$\pm 0.015m$, the Y-axis within ±$\pm 0.003m$, and the Z-axis within ±$\pm 0.2m$. Likewise, the oval (cross-medium) trajectory following results, as shown in Fig. 12, also demonstrates the effectiveness of our controller.
V-C Reinforcement Learning from Demonstration Evaluation
Fig. 13 demonstrates successful training of RLfD in our SEALS simulator for our AAM robot.
Figure 13: Reinforcement Learning from Demonstration results
V-D Visual Reinforcement Learning Evaluation
We applied the Soft Actor-Critic algorithm [11], a robust reinforcement learning (RL) method for continuous control, to train our AAM in the SEALS environment to reach objects such as crabs in Fig. 8, using distance-based rewards. The observation space included two 128 x 128 RGB images from the front and bottom cameras on the AAM (resulting in six channels in total) and the global poses of the AAM and the target object. The cumulative average rewards over the past 100 episodes, shown in Fig. 8, indicate that the reinforcement learning converges. The learning results is shown in Fig. 14. For more detailed robot learning results, please refer to Appendix C.
@@ VI CONCLUSION
Our work makes significant contributions to the field of robotics by introducing a new class of robots, Aerial-Aquatic Manipulators (AAMs), and developing the first high-fidelity simulator that integrates sea, air, and land environments. The unique benefits of this work include enabling AAMs to seamlessly transition between different environments and perform complex tasks that require cross-medium manipulation.
Limitations: While our research presents a robust simulation environment, we acknowledge the limitation of not being able to fully verify the Sim2Real transfer. Developing a fully functional physical AAM involves substantial research efforts, such as complex electronics design, achieving IP68 waterproofing, lightweight structures, and battery safety, that could constitute a separate study. In essence, building physical and simulated AAMs presents a chicken-and-egg problem, where each development phase influences the other.
Nonetheless, AAM-SEALS opens a broad avenue for future research efforts:
  • Supporting the Development of Physical AAMs: Leveraging AAM-SEALS to aid in the development of physical AAMs is an ongoing project. This simulator will provide critical insights and validation before constructing physical prototypes.
  • Modeling Power Depletion Effects: Future versions of AAM-SEALS could incorporate the impact of battery depletion on motors, control, and planning, as exemplified in [6]. This addition would enhance the realism and practical utility of the simulations, allowing for more accurate modeling of how power limitations affect AAM performance.
@@
  • Efficient Simulation: Simulating particle physics is computationally expensive. Future research could focus on developing a hierarchical simulation approach, where high-resolution simulations are limited to the local region around the AAM. This localized high-resolution simulation would be informed by an outer, lower-resolution simulation that incorporates broader environmental factors such as temperature, depth, and ocean flow.
@@ Appendix A Potential Questions and Our Answers
Not necessarily. Many advanced simulator studies (e.g., [37, 41, 20, 10, 4, 35]) prioritize simulation development over physical experiments due to practical constraints. Building on these works, we validated our simulator’s hydrodynamic accuracy through real-world experiments with 3D-printed AAM models, demonstrating realistic acceleration patterns during air-to-water transitions. While these tests confirm hydrodynamic accuracy, they do not fully capture the complexity of cross-medium tasks. This underscores the role of AAM-SEALS as a crucial foundation for guiding future physical AAM development.
@@
It’s great that you have evaluated Reinforcement Learning from Demonstrations techniques, such as Soft Actor-Critic from Demonstrations (SACfD). Did you also consider standard Reinforcement Learning methods, like Soft Actor-Critic (SAC)?
Yes, we have evaluated SAC as well. While it is effective, it performs slightly worse than SACfD – it requires more training episodes to achieve meaningful performance improvements and ultimately converges to a lower score. Please refer to Fig. 15 for detailed results.
Figure 15: Reinforcement Learning (Soft-Actor Critic) results
@@ Appendix C Additional Reinforcement Learning Experiments
C-A Reinforcement Learning from Demonstrations Hyperparameters
@@
replay buffer size 100,000
@@
N-Step 10
pretrain step 5,000
PER_Alpha 0.6
PER_Beta 0.4
PER_EPS 1e-6
PER_EPS_DEMO 1.0